OpenStudy (anonymous):

write the slope-intercept form of the equation of the the described. through (-3,5), parallel to y=-8/3x+1

OpenStudy (kc_kennylau):

\[\Large y=-\frac83x+1\]

OpenStudy (kc_kennylau):

So, what do two parallel lines have in common?

OpenStudy (anonymous):

they never meet?

OpenStudy (kc_kennylau):

Their slope is the same xd

OpenStudy (kc_kennylau):

What is the slope of the second line?

OpenStudy (anonymous):

y=mx+b 5=-8/3(-3) -24/3=-8

OpenStudy (kc_kennylau):

y=mx+b 5=(-8/3)(-3)+b 5=-8+b 13=b b=13

OpenStudy (anonymous):

thanks!!